### v5

@@ Abstract
DNA sequence alignment involves assigning short DNA reads to the most probable locations on an extensive reference genome. This process is crucial for various genomic analyses, including variant calling, transcriptomics, and epigenomics. Conventional methods, refined over decades, tackle this challenge in two steps: genome indexing followed by efficient search to locate likely positions for given reads. Building on the success of Large Language Models (LLM) in encoding text into embeddings, where the distance metric captures semantic similarity, recent efforts have explored whether the same Transformer architecture can produce numerical representations for DNA sequences. Such models have shown early promise in tasks involving classification of short DNA sequences, such as the detection of coding- vs non-coding regions, as well as the identification of enhancer and promoter sequences. Performance at sequence classification tasks does not, however, translate to sequence alignment, where it is necessary to conduct a genome-wide search to align every read successfully, a significantly longer-range task by comparison. We bridge this gap by framing the Sequence Alignment task for Transformer models as an “Embed-Search-Align” task. In this framework, a novel Reference-Free DNA Embedding RDE model generates representations of reads and fragments of the reference, which are projected into a shared vector space where the read-fragment distance is used as a surrogate for alignment. In particular, we introduce: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich reference-free, sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale. RDE is 99% accurate when aligning 250-length reads onto a human reference genome of 3 gigabases (single-haploid), rivaling conventional algorithmic sequence alignment methods such as Bowtie and BWA-Mem. RDE far exceeds the performance of $6$ recent DNA-Transformer model baselines on ESA such as Nucleotide Transformer, Hyena-DNA, and shows task transfer across chromosomes and species.
@@
Sequence alignment is a central problem in the analysis of sequence data. It is used in various genomic analyses, including variant calling, transcriptomics, and epigenomics. Many DNA sequencers generate short reads that are only a couple of hundred bases long. In order to interpret this data, a typical first step is to align the reads to a genome. Genomes come in many sizes, but many of the most commonly studied genomes, such as that of humans, are three gigabases long. Moreover, most experiments will generate millions of reads. The resulting task of aligning short reads to large-size genomes is computationally challenging –akin to finding a pin in a haystack– and many decades of work have led to optimized approaches that can perform these tasks with great efficiency.
Since genomes are sequences where the alphabet consists of only a few symbols ($\{A,T,G,C\}$) they can be considered as Limited Alphabet Languages (LAL). We ask whether one can design a new paradigm to align reads to genomes by exploiting the Transformer architectures vaswani_attention_2017 adopted for recent language modeling efforts. The applications of the Transformer model – and, more generally “Large Language Models” (LLM) – in Bioinformatics applications are still in their infancy, yet hold substantial promise: Transformer models have demonstrated an unprecedented ability to construct powerful numerical representations of sequential data vit ; wav2vec ; gpt3 .
Numerous DNA Transformer models (ji_dnabert_2021, ; zvyagin_genslms_2022, ; fishman_gena-lm_2023, ; nucleotide, ) have emerged recently, mainly designed for classification tasks in downstream applications. However, these models do not explicitly consider a fundamental distinction between Limited Alphabet Languages (LAL) (such as genomes) and natural languages with large vocabulary. In LALs, there is a one-to-one correspondence between the precise symbol ordering and the underlying “meaning”. If a protein is encoded by one amino acid sequence, a slightly different sequence (with a few edits) would encode for a different protein with a distinct functionality in the cells. These models, trained on classification tasks, generate sequence embeddings such that their pairwise distances correspond to class separation. As a result, pairs of sequences with very large edit distances between them are mapped to numerical representations that are close by. However, in tasks such as Sequence Alignment the objective is quite different: the pairwise representation distance has to closely match the sequence edit distance. Indeed as portrayed in Figure 1, current DNA Transformer models without any further refinement are unable to perform Sequence Alignment tasks: the representation space is simply not designed for such an objective. This demonstrates the need for a new model which is able to transfer a precise and computationally expensive distance metric, such as the edit distance, over LALs to a continuous representation space.
@@
We establish a foundation model foundation Reference-free DNA Embedding (RDE) – tailored for embedding DNA sequences. The model $h_{\Theta}$ (i.e., parameterized by weights $\Theta$) maps any genome subsequence $F$ into an embedding $h_{\Theta}(F)\in\mathcal{R}^{n}$ (e.g., $n=768)$, such that if $F_{1}$ and $F_{2}$ are similar subsequences of differing lengths – for example, $F_{1}$ is any noisy subsequence of $F_{2}$ or $F_{1}$ and $F_{2}$ have significant overlaps – then the $d(h_{\Theta}(F_{1}),h_{\Theta}(F_{2}))$ is small; where $d(\cdot)$ is a distance metric in the embedding space, such as the cosine distance between two vectors. Another requirement of such RDE models is that they should generate embeddings that are reference free, i.e. once trained, it creates semantically-aware embedding (i.e. sequences that are close in edit distance are mapped close to each other) of any given genome subsequences irrespective of the reference genome from which it is sampled. Given such a model, alignment of reads can be achieved by a near-neighbor search in the embedding space: only the fragments of genome with embeddings nearest to the embedding of the given read need to be considered. Thus, a global search in a giga-bases long sequence is reduced to a local search in a vector space.
$r:=(\tilde{b}_{1},\tilde{b}_{2},\dots,\tilde{b}_{Q}),$ (1)
[…]
such that the edit distance $d(r,\tilde{r})$ – for example the Smith-Waterman (SW) distance – is minimized. The primary objective is to identify the most probable location, $q$, of this read within the reference genome.
@@
Computational simulators have been developed to generate synthetic reads that have properties of real reads. These simulators mimic the read quality and characteristics produced by actual sequencing machines, thus providing a scalable means for validating new alignment approaches (huang_art_2012, ).
@@
Conventional sequence alignment methods, such as those used in BWA-MEM, have evolved significantly from the basic SW approach, which computes the similarity between a read $r$ and reference $\mathcal{R}$; using the SW distance on its own, however, is computationally intensive for large genomes. Some key advances include: (a) Sharding the reference into smaller fragments for individual searching; (b) Progressive search using phylogenetic trees and distance heuristics for logarithmic complexity scaling; (c) Compression techniques like the Burrows-Wheeler transform (bw_transform, ), reducing the effective search length; and (d) Multi-core/thread implementations and database instantiations for faster computations and improved data recall (minimap, ; bwamem2, ; bowtie, ).
In this work, we explore an alternative paradigm for aligning a read to a genome, drawing parallels with advancements in Natural Language Processing (NLP). Traditional NLP relied on rule-based techniques to diagram sentences into grammatical components, using dependency parse trees to discern semantic relationships. However, the introduction of models like the Transformers (vaswani_attention_2017, ) revolutionized this approach. Unlike rule-based frameworks, Transformers implicitly identify syntactic and semantic structures tenney , as evidenced by their performance at various NLP tasks like Sentiment Analysis, Entity Recognition, and Question-Answering (devlin_bert_2019, ; reimers_sentence-bert_2019, ; reviewllm, ). The flexibility of this architecture has facilitated its application beyond language processing, including in vision (dosovitskiy_image_2020), auditory (audiotransformer, ), and neurological domains (wang_brainbert_2023, ). In bioinformatics, researchers are now harnessing Transformers to bypass genome-specific hard-coded rules, a development we contribute to with our novel method for Sequence Alignment.
@@
DNA sequences share remarkable similarities with written language, offering a compelling avenue for the application of Transformer models. Like written language, these are sequences generated by a small alphabet of nucleotides $\{A,T,G,C\}$. Traditional DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (huffman_method_1952, ; manber_suffix_1993, ) – to successfully parse and compress DNA sequences.
Recognizing such similarities, within the last few years, several Transformer-based models have been developed for DNA sequence analysis. Notably, DNABERT-2 (ji_dnabert_2021, ; dna2bert, ), Nucleotide Transformer (nucleotide, ), GenSLM (zvyagin_genslms_2022, ), HyenaDNA (nguyen_hyenadna_2023, ) and GENA-LM (fishman_gena-lm_2023, ) have been designed to discern relationships between short genetic fragments and their functions. Specifically, Nucleotide Transformer representations have shown utility in classifying key genomic features such as enhancer regions and promoter sequences. Similarly, GENA-LM has proven effective in identifying enhancers and Poly-adenylation sites in Drosophila. In parallel, DNABERT-2 representations have also been found to cluster in the representation space according to certain types of genetic function.
@@
In this paper, we argue that both limitations L1, L2 of Transformer-DNA models can be mitigated by formulating sequence alignment as a vector search-and-retrieval task. Our approach is twofold: (A) We introduce a sequence encoder Reference-Free DNA Embedding (RDE), trained through self-supervision, to embed DNA subsequences to vectors. (B) Next, we formulate a framework, Embed-Search-Align (ESA), that maps reads to reference genome sequence. We leverage a specialized data structure, termed a DNA vector store, that enables efficient storage and search in an embedding space [^2] Codebase available here: https://anonymous.4open.science/r/dna2vec-7E4E/: the entire reference genome is sharded into equal-length overlapping fragments whose embeddings are then uploaded to the vector store. For each read, top-K closest fragments can then be searched efficiently in time that scales only logarithimically in the number of fragments. These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance – over explicit pooling methods – at abstractive semantic tasks such as evidence retrieval lewis and semantic text similarity (gao_simcse_2021, ; chen_simple_2020, ). (B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (johnson_billion-scale_2019, ) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
2.1 2.1 RDE: Designing effective sequence representations
An optimal sequence encoder model $h$ is such that the corresponding embeddings of any read $r$ and reference fragment $\mathcal{F}$$h(r),h(\mathcal{F})$ respectively – obey the following constraints over a pre-determined distance metric $d$:
2.2 2.2 RDE: Self-supervision and contrastive loss
2.3 2.3 RDE: Encoder implementation
RDE uses a Transformer-encoder (devlin_bert_2019, ; vaswani_attention_2017, ), comprising $12$ heads and $6$-layers of encoder blocks. The size of the vocabulary is $10,000$. Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps. Generated numerical representations for each read (and fragment) are projected into a high-dimensional vector space, $h(r)\in\mathbb{R}^{1020}$. The learning rate is annealed using one-cycle cosine annealing (smith_super-convergence_2019, ), dropout is set to $0.1$, and $\tau=0.05$. During training, the reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}([800,2000])$ and read $|r_{i}|\sim\mathcal{U}([150,500])$ have variable lengths sampled from a uniform distribution. Here, $|x|$ denotes the length of $x$. To improve model performance on noisy reads, $1-5\%$ of bases are replaced with another random base in $40\%$ of the reads in a batch. Shorter sequences were padded to equal the length of the longest sequence in a batch. The distance metric used is Cosine Similarity.
2.4 2.4 ESA: Task Definition
@@
Indexing: For a given reference genome $\mathcal{R}$, we construct a minimal set of reference fragments :=$\mathcal{F}:=\{\mathcal{F}_{1},\mathcal{F}_{2},\dots\}$ to span $\mathcal{R}$. Note that the fragments overlap at least a read length; i.e. $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|\geq Q$ to guarantee that every read is fully contained within some fragment in the set. In our experiments with external read generators (huang_art_2012, ), $Q_{max}=250,|\mathcal{F}_{i}|=1250$. Each reference fragment is encoded using the trained RDE model, and the resulting sequence embeddings ($\in\mathbb{R}^{1020}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database. Once populated with all the fragments, we are ready to perform the alignment.
@@
The results from Sec. 3 demonstrate that even for pure reads, baseline models do not generate adequate representations to perform sequence alignment. In this section, RDE and the two best baselines – DB2, mean and HD, max – are evaluated on ESA using reads generated from an external read simulator (ART) – see huang_art_2012 . ART has served as a reliable benchmark for evaluating other contemporary alignment tools and provides controls to model mutations and variations common in reads generated by Illumina machines.
@@
Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve (6), and for this, we need to compute the SW score. For all presented results, the settings are: match_score = -2, mismatch_penalty = +1, open_gap_penalty = +0.5, continue_gap_penalty = +0.1. After alignment, we get $q^{*}$ – see (7) – as the estimated location of a read in the genome. Let $\hat{q}^{*}$ be its true location. If $q^{*}=\hat{q}^{*}$, it is a perfect match and the recall is successful. In cases where there is a mutation in the first or last position in a read, the fine-alignment will return $q^{*}_{\mathcal{F}^{*}}$ offset by at most $2$ locations, resulting in $\hat{q}^{*}=q^{*}\pm 2$. Hence, the condition for an exact location match: $|q^{*}-\hat{q}^{*}|\leq 2$.
@@
While our first-generation reference-free DNA embedding (RDE) models provide alignment performance that almost matches the performance of traditional aligners (refined over decades), there is room for considerable improvement. First, our RDE models have been trained using samples that span only around 2% of the human genome as shown in the Sec-A of the Supporting Material(SM). One needs to explore if the embedding properties of our models can be further improved by diversifying and augmenting training data and optimizing over the model parameters. The validation of RDE models can be performed only through various genomic tasks. In terms of the task of alignment, we plan to explore the following: (i) The current off-the-shelf implementation has a speed of around 10K reads per minute; while the existing aligners can achieve close to 1M reads per minute by trading off alignment accuracy. As detailed in SM Sec. B, we are considering various optimization strategies, including model compilation, to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment. We believe ESA can be made much faster; (ii) Optimization of the training of the RDE models to improve performance for shorter reads. As shown in see Table 2, Page 3 of SM the existing model performs well on longer reads; we want to develop RDE models that specialize to aligning shorter reads; and (iii) quantifying how embeddings change as subsequences are edited; this will help in better alignment accuracy. It will also help in correlating the geometry of the embedding space to the structure of DNA sequences.
@@
As a first validation and application of RDE, we developed an Embed-Align-Search (ESA) framework where alignment of reads to a reference genome is achieved by a near-neighbor search in the embedding space: only the fragments of reference genome with embeddings nearest to the embedding of the given read need to be considered. Thus, a global search in a giga-bases long reference genome sequence is reduced to a local search in a vector space. Detailed experiments showed that the performance of even this first-generation RDE methodology is comparable to that of traditional aligners, which have been refined over decades. Future differentiating application of RDE to the alignment task might lie in aligning reads to graph- or pan-genomes liao , where the reference genome has multiple variants. Indexing such pangenomes is a difficult task in traditional aligners; however, in ESA one can store the RDE embeddings of the different fragments in the pangenome, and then perform a near-neighbor search in the embedding space. Thus, the ESA methodology does not have to be significantly altered, and alignment of reads to pangenomes constitutes an important future task.
Going forward we aim to answer the following important questions:
  • How reference-free is our RDE model? To benchmark the generalizability of our RDE model we have performed several cross-species alignment tasks as summarized in Tables 3 and 4 in the SM. In particular, we have used an RDE trained using human genome to align reads to reference genomes belonging to a wide range of species. The results are encouraging and the design of universal RDEs –that would generate effective embeddings of subsequences sampled from the genome of every species – is another direction of future research.
  • Is the RDE capable of generalizing to other tasks like de novo Genome Assembly? One can also harness the RDE space to address the complementary task of de novo Genome Assembly, where given a set of reads one assembles a genome by stitching together the reads, without using any reference. Genome assembly is considered to be a much more challenging task than alignment. In SM Sec. E, we provide initial proofpoints of a genome assembly framework, under reasonable simplifying assumptions about read lengths and overlaps among the reads. We show how the RDE model – trained on exemplars from human genome – can effectively embed simulated reads from a very different species, Thermus Aquaticus. We find that the read embeddings form a low-dimensional manifold, which can be efficiently searched locally to obtain high-fidelity genome assembly from the simulated reads. Our future work on genome assembly will focus on improving RDE models to obtain better embeddings, and also on developing local-search based algorithms that can accurately map low-dimensional manifolds (comprised by the embeddings of reads) into linear DNA sequences.

### v6

@@ Abstract
DNA sequence alignment, an important genomic task, involves assigning short DNA reads to the most probable locations on an extensive reference genome. Conventional methods tackle this challenge in two steps: genome indexing followed by efficient search to locate likely positions for given reads. Building on the success of Large Language Models (LLM) in encoding text into embeddings, where the distance metric captures semantic similarity, recent efforts have encoded DNA sequences into vectors using Transformers and have shown promising results in tasks involving classification of short DNA sequences. Performance at sequence classification tasks does not, however, guarantee sequence alignment, where it is necessary to conduct a genome-wide search to align every read successfully, a significantly longer-range task by comparison. We bridge this gap by developing a “Embed-Search-Align” (ESA) framework, where a novel Reference-Free DNA Embedding (RDE) Transformer model generates vector embeddings of reads and fragments of the reference in a shared vector space; read-fragment distance metric is then used as a surrogate for sequence similarity. ESA introduces: (1) Contrastive loss for self-supervised training of DNA sequence representations, facilitating rich reference-free, sequence-level embeddings, and (2) a DNA vector store to enable search across fragments on a global scale. RDE is 99% accurate when aligning 250-length reads onto a human reference genome of 3 gigabases (single-haploid), rivaling conventional algorithmic sequence alignment methods such as Bowtie and BWA-Mem. RDE far exceeds the performance of $6$ recent DNA-Transformer model baselines such as Nucleotide Transformer, Hyena-DNA, and shows task transfer across chromosomes and species.
@@
Sequence alignment is a central problem in the analysis of sequence data. It is used in various genomic analyses, including variant calling, transcriptomics, and epigenomics. Many DNA sequencers generate short reads that are only a couple of hundred bases long. In order to interpret this data, a typical first step is to align the reads to a genome. Genomes come in many sizes, and commonly studied genomes range from millions of bases (e.g., bacteria) to billions of bases (e.g., mammals, plants) in length. The resulting task of aligning short reads to large-size genomes is computationally challenging—akin to finding a needle in a haystack—and many decades of work have led to optimized approaches that can perform these tasks with great efficiency. Since genomes are sequences where the alphabet consists of only a few symbols ($\{A,T,G,C\}$) they can be considered as Limited Alphabet Languages (LAL). We ask whether one can design a new paradigm to align reads to genomes by exploiting the Transformer architectures vaswani_attention_2017 adopted for recent language modeling efforts. The applications of the Transformer model—and, more generally “Large Language Models” (LLM)—in Bioinformatics applications are still in their infancy, yet hold substantial promise: Transformer models have demonstrated an unprecedented ability to construct powerful numerical representations of sequential data vit ; wav2vec ; gpt3 .
@@
We establish a foundation model foundation –Reference-free DNA Embedding (RDE)—tailored for embedding DNA sequences. The model $h_{\Theta}$ (i.e., parameterized by weights $\Theta$) maps any genome subsequence $F$ into an embedding $h_{\Theta}(F)\in\mathcal{R}^{n}$ (e.g., $n=768)$, such that if $F_{1}$ and $F_{2}$ are similar subsequences of differing lengths—for example, $F_{1}$ is any noisy subsequence of $F_{2}$ or $F_{1}$ and $F_{2}$ have significant overlaps—then the $d(h_{\Theta}(F_{1}),h_{\Theta}(F_{2}))$ is small; where $d(\cdot)$ is a distance metric in the embedding space, such as the cosine distance between two vectors. Another requirement of such RDE models is that they should generate embeddings that are reference free, i.e. once trained, it creates semantically-aware embedding (i.e. sequences that are close in edit distance are mapped close to each other) of any given genome subsequences irrespective of the reference genome from which it is sampled. Given such a model, alignment of reads can be achieved by a near-neighbor search in the embedding space: only the fragments of genome with embeddings nearest to the embedding of the given read need to be considered. Thus, a global search in a giga-bases long sequence is reduced to a local search in a vector space.
@@
DNA sequences share remarkable similarities with written language, offering a compelling avenue for the application of Transformer models. Like written language, these are sequences generated by a small alphabet of nucleotides $\{A,T,G,C\}$. Traditional DNA modeling efforts have already accommodated mature encoding and hashing techniques initially developed for written language – such as Suffix trees/arrays and Huffman coding (huffman_method_1952, ; manber_suffix_1993, ) – to successfully parse and compress DNA sequences. Alignment methods such as MinHash and MashMap mashmap have incorporated Locality Sensitive Hashing (LSH) lsh to construct fast, approximate representations of DNA sequences that facilitate the rapid matching of millions of reads onto the reference genome.
Within the last few years, several Transformer-based models have been developed for DNA sequence analysis. Notably, DNABERT-2 (ji_dnabert_2021, ; dna2bert, ), Nucleotide Transformer (nucleotide, ), GenSLM (zvyagin_genslms_2022, ), HyenaDNA (nguyen_hyenadna_2023, ) and GENA-LM (fishman_gena-lm_2023, ) have been designed to discern relationships between short genetic fragments and their functions. Specifically, Nucleotide Transformer representations have shown utility in classifying key genomic features such as enhancer regions and promoter sequences. Similarly, GENA-LM has proven effective in identifying enhancers and Poly-adenylation sites in Drosophila. In parallel, DNABERT-2 representations have also been found to cluster in the representation space according to certain types of genetic function.
@@
In this paper, we argue that both limitations L1, L2 of Transformer-DNA models can be mitigated by formulating sequence alignment as a vector search-and-retrieval task. Our approach is twofold: (A) We introduce a sequence encoder Reference-Free DNA Embedding (RDE), trained through self-supervision, to embed DNA subsequences to vectors. (B) Next, we formulate a framework, Embed-Search-Align (ESA), that maps reads to reference genome sequence. We leverage a specialized data structure, termed a DNA vector store, that enables efficient storage and search in an embedding space [^1] Codebase available here: https://anonymous.4open.science/r/dna2vec-7E4E/: the entire reference genome is sharded into equal-length overlapping fragments whose embeddings are then uploaded to the vector store. For each read, top-K closest fragments can then be searched efficiently in time that scales only logarithmically in the number of fragments hnsw . These strategies have been explored in NLP: (A) Sequence-to-embedding training using contrastive loss has shown improved performance—over explicit pooling methods—at abstractive semantic tasks such as evidence retrieval lewis and semantic text similarity (gao_simcse_2021, ; chen_simple_2020, ). (B) Specialized data structures, such as “vector stores” or “vector databases” like FAISS (johnson_billion-scale_2019, ) and Pinecone, use advanced indexing and retrieval algorithms for scalable numerical representation search.
2.1 2.1 Task Definition
$r:=(\tilde{b}_{1},\tilde{b}_{2},\dots,\tilde{b}_{Q}),$ (1)
[…]
such that the edit distance $d(r,\tilde{r})$—computed using the Smith-Waterman (SW) algorithm—is minimized. The primary objective is to identify the most probable location, $q$, of this read within the reference genome.
@@
2.2 2.2 RDE: Designing effective sequence representations
An optimal sequence encoder model $h$ is such that the corresponding embeddings of any read $r$ and reference fragment $\mathcal{F}$$h(r),h(\mathcal{F})$ respectively—obey the following constraints over a pre-determined distance metric $d$:
2.3 2.3 RDE: Self-supervision and contrastive loss
2.4 2.4 RDE: Encoder implementation
RDE uses a Transformer-encoder (devlin_bert_2019, ; vaswani_attention_2017, ), comprising $12$ heads and $6$-layers of encoder blocks. The size of the vocabulary is $10,000$. Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps. Generated numerical representations for each read (and fragment) are projected into a high-dimensional vector space, $h(r)\in\mathbb{R}^{1020}$. The learning rate is annealed using one-cycle cosine annealing (smith_super-convergence_2019, ), dropout is set to $0.1$, and $\tau=0.05$. During training, the reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}([800,2000])$ and read $|r_{i}|\sim\mathcal{U}([150,500])$ have variable lengths sampled from a uniform distribution. Here, $|x|$ denotes the length of $x$. To improve model performance on noisy reads, $1-5\%$ of bases are replaced with another random base in $40\%$ of the reads in a batch. Shorter sequences were padded to equal the length of the longest sequence in a batch. The similarity measure used is Cosine Similarity.
@@
Indexing: For a given reference genome $\mathcal{R}$, we construct a minimal set of reference fragments :=$\mathcal{F}:=\{\mathcal{F}_{1},\mathcal{F}_{2},\dots\}$ to span $\mathcal{R}$. Note that the fragments overlap at least a read length; i.e. $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|\geq Q$ to guarantee that every read is fully contained within some fragment in the set. In our experiments with external read generators (huang_art_2012, ), $Q_{max}=250,|\mathcal{F}_{i}|=1250$. Each reference fragment is encoded using the trained RDE model, and the resulting sequence embeddings ($\in\mathbb{R}^{1020}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone [^3] https://www.pinecone.io database. Once populated with all the fragments, we are ready to perform the alignment.
@@
The results from Sec. 3 demonstrate that even for pure reads, baseline models do not generate adequate representations to perform sequence alignment. In this section, RDE and the two best baselines—DB2, mean and HD, max—are evaluated on ESA using reads generated from an external read simulator (ART) – see huang_art_2012 . ART has served as a reliable benchmark for evaluating other contemporary alignment tools and provides controls to model mutations and variations common in reads generated by Illumina machines.
@@
Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve (4), and for this, we need to compute the SW score. For all presented results, the settings are: match_score = -2, mismatch_penalty = +1, open_gap_penalty = +0.5, continue_gap_penalty = +0.1. After alignment, we get $q^{*}$—see (7)—as the estimated location of a read in the genome. Let $\hat{q}^{*}$ be its true location. If $q^{*}=\hat{q}^{*}$, it is a perfect match and the recall is successful. In cases where there is a mutation in the first or last position in a read, the fine-alignment will return $q^{*}_{\mathcal{F}^{*}}$ offset by at most $2$ locations, resulting in $\hat{q}^{*}=q^{*}\pm 2$. Hence, the condition for an exact location match: $|q^{*}-\hat{q}^{*}|\leq 2$.
@@
While our first-generation reference-free DNA embedding (RDE) models provide alignment performance that almost matches the performance of traditional aligners (refined over decades), there is room for considerable improvement. First, our RDE models have been trained using samples that span only around 2% of the human genome as shown in the Sec-A of the Supporting Material(SM). One needs to explore if the embedding properties of our models can be further improved by diversifying and augmenting training data and optimizing over the model parameters. The validation of RDE models can be performed only through various genomic tasks. In terms of the task of alignment, we plan to explore the following: (i) The current off-the-shelf implementation of our ESA algorithm has a speed of around 10K reads per minute. The existing aligners such as Bowtie are faster, and can achieve close to 1M reads per minute; additional speedups can be achieved by trading off alignment accuracy vargas . As detailed in SM Sec. B, we are considering various optimization strategies, including model compilation, to speed up inference and enhanced parallelization in vector store searches and fragment-read alignment. We believe ESA can be made much faster; (ii) Optimization of the training of the RDE models to improve performance for shorter reads. As shown in see Table 2, Page 3 of SM the existing model performs well on longer reads; we want to develop RDE models that specialize to aligning shorter reads; and (iii) quantifying how embeddings change as subsequences are edited; this will help in better alignment accuracy. It will also help in correlating the geometry of the embedding space to the structure of DNA sequences.
@@
As a first validation and application of RDE, we developed an Embed-Align-Search (ESA) framework where alignment of reads to a reference genome is achieved by a near-neighbor search in the embedding space: only the fragments of reference genome with embeddings nearest to the embedding of the given read need to be considered. Thus, a global search in a giga-bases long reference genome sequence is reduced to a local search in a vector space. Detailed experiments showed that the performance of even this first-generation RDE methodology is comparable to that of traditional aligners, which have been refined over decades. Future application of RDE to the alignment task might lie in aligning reads to pan-genomes liao , where the diversity in the genomic content – modeled as allelic paths in a reference genome graph – can be exhaustively mapped onto the described embedding space (as in Fig.  2). Conveniently, the proposed nearest-neighbor search would align a given read to fragments; the only difference being that multiple fragments can belong to the same reference location but different reference paths. The ESA methodology would not need to be modified beyond this.
An important question we want to answer going forward is: How reference-free is our RDE model? To benchmark the generalizability of our RDE model we have performed several cross-species alignment tasks as summarized in Tables 3 and 4 in the SM. In particular, we have used an RDE trained using human genome to align reads to reference genomes belonging to a wide range of species. The results are encouraging and the design of universal RDEs—that would generate effective embeddings of subsequences sampled from the genome of every species—is another direction of future research.
One can also harness the RDE space to address the complementary task of de novo Genome Assembly, where given a set of reads one assembles a genome by stitching together the reads, without using any reference. Genome assembly is considered to be a much more challenging task than alignment. In SM Sec. F, we provide initial proofpoints of a genome assembly framework, under reasonable simplifying assumptions about read lengths and overlaps among the reads. We show how the RDE model—trained on exemplars from human genome—can effectively embed simulated reads from a very different species, Thermus Aquaticus. We find that the read embeddings form a low-dimensional manifold, which can be efficiently searched locally to obtain high-fidelity genome assembly from the simulated reads. Our future work on genome assembly will focus on improving RDE models to obtain better embeddings, and also on developing local-search based algorithms that can accurately map low-dimensional manifolds (comprised by the embeddings of reads) into linear DNA sequences.